# Physics-informed DeepCT: Sinogram Wavelet Decomposition Meets Masked Diffusion

### v2

@@ III-C Cascade Reconstruction of Sinogram and Wavelet
where $dt$ is an infinitesimal negative time step, $\bar{w}$ is a standard Brownian motion with time flows backwards from $T\rightarrow 0$. $\nabla_{\hat{y}}\log p_{t}(\hat{y})$ is the score for each marginal distribution, $\hat{y}=\{\hat{y}_{s},\hat{y}_{h}\}$ includes the sinogram and its wavelet high-frequency regions. $\nabla_{\hat{y}}\log p_{t}(\hat{y})$ can be estimated by training a time-dependent scoring network $\textbf{s}_{\theta}(\hat{y},t)=\{\textbf{s}_{\theta}(\hat{y}_{s},t),\textbf{s}_{\theta}(\hat{y}_{h},t)\}$ to satisfy $\textbf{s}_{\theta}(\hat{y},t)\simeq\nabla_{\hat{y}}\log p_{t}(\hat{y})$. The score function $\theta^{\ast}$ can be estimated by network training in the learning stage of prior information. The scoring estimator $\textbf{s}_{\theta}(\hat{y}_{s,t},t)$ can replace Eq. (15) to approximate the solution of the score function:
@@
where $\tilde{H}[\cdot]$ represents the full-frequency information, $\tilde{H}_{l}[\cdot]$ represents the low-frequency component and $\tilde{H}_{h}[\cdot]$ represents the high-frequency component. Through the iterative optimization mechanism of SRM and SHD, a smooth transition from global consistency to detail precision is effectively achieved.
1) Sinogram Generation: The sinogram generation stage enhances the global structural information in the image, thereby ensuring the consistency and integrity of the data during the processing. In this stage, it can be further refined into two interrelated subtasks, which respectively focus on the modeling of global information and the expression and optimization of structural features, so as to improve the expressive ability of the overall data and the reconstruction accuracy. This stage can be decomposed into the following two sub-problems:
@@
$\hat{y}_{s}^{t-1}=\underset{\hat{y}_{s}}{\arg\min}\ \|\hat{y}_{s}-\hat{y}_{s}^{t-\frac{1}{2}}\|^{2}_{2}+\nu_{s}R_{s}(\hat{y}_{s}).\vspace{-0.5 em}$ (20)
@@
$\hat{y}_{s}=(1-P(\wedge))\hat{y}_{s}+P(\wedge)y.\vspace{-0.5 em}$ (21)
@@
2) High-frequency generation in wavelet domain: After the preliminary reconstruction is completed in the projection domain during the first stage, the obtained sinogram is mapped to the wavelet domain. In this domain, the image is decomposed at multiple scales, from which the high-frequency components in three directions are extracted, and the low-frequency components are retained simultaneously, so as to fully capture the detailed structures and the overall contours in the image. The process can be described as follows:
@@
3) Domain transform stage: Finally, this process integrates the previously retained low-frequency information with the optimized high-frequency components. By fusing information across different frequency levels, a more comprehensive image representation is achieved. Subsequently, the combined data is mapped back to the projection domain through the inverse wavelet transform, as detailed below:
@@
$\tilde{x}=FBP(\hat{y}).$ (29)
In summary, the iterative reconstruction stage of SWARM is shown in Fig. 2(c). By embedding a random mask and introducing a random high-frequency strategy in the sinogram, the prior distributions of global structural and detailed features are extracted, effectively guiding the training process of the model. In the actual reconstruction, through the iterative updates of the numerical SDE solver and the Langevin dynamics, high-quality full-view projection data is gradually obtained. The specific reconstruction process is shown in Fig. 2.
Alg. 1 provides a detailed description of SRM and SHD in the training process. The detailed process of the reconstruction stage is shown in Alg. 2.
Algorithm 1 Training process.
@@
   SRM Training process:
   1:  Sample a projected data $y$;
   2:  Randomly generated a discrete random mask $\hat{m}$;
   3:  Generate a masked sinogram: $\tilde{Y}_{s}=y\odot\hat{m}$;
   4:  Trained with SRM;
   5:  return $\textbf{s}_{\theta_{s}}(\tilde{Y}_{s},t)$.
   SHD Training process:
   1:  Sample a projected data $y$;
   2:  Wavelet transform:
   3:  $H:y\rightarrow\{h_{LL}(y),h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$;
   4:  Random selection of high-frequency components:
   5:  $\tilde{Y}_{h}\leftarrow\{h_{LH}(y),h_{HL}(y),h_{HH}(y)\}$;
   6:  Trained with SHD;
   7:  return $\textbf{s}_{\theta_{h}}(\tilde{Y}_{h},t)$.
Algorithm 2 Iterative Reconstruction process.
@@
   1:   $\hat{Y}_{s}^{T}\sim\mathcal{N}(0,\sigma^{2}_{max})$ , $\hat{Y}_{h}^{T}\sim\mathcal{N}(0,\sigma^{2}_{max})$;
   2:  For $t=T-1$ to 00 do:
   3:     $\hat{Y}_{s}^{t-\frac{1}{2}}\leftarrow predictor(\hat{Y}_{s}^{t},\sigma_{t-1},\sigma_{t},\textbf{s}_{\theta_{s}})$;
   4:     Update $\hat{Y}_{s}^{t-\frac{1}{2}}$ with Eq. (21);
   5:     $\hat{Y}_{s}^{t-1}\leftarrow corrector(\hat{Y}_{s}^{t-\frac{1}{2}},\sigma_{t-1},\varepsilon_{t-1},\textbf{s}_{\theta_{s}})$;
   6:     Update $\hat{Y}_{s}^{t-1}$ with Eq. (21);
   7:     For $i=1$ to $3$ do:
   8:        $\hat{Y}_{h,i}^{t-\frac{1}{2}}\leftarrow predictor(\hat{Y}_{h,i}^{t},\sigma_{t-1},\sigma_{t},\textbf{s}_{\theta_{h}})$;
   9:        Update $\hat{Y}_{h,i}^{t-\frac{1}{2}}$ with Eq. (25);
   10:       $\hat{Y}_{h,i}^{t-1}\leftarrow corrector(\hat{Y}_{h,i}^{t-\frac{1}{2}},\sigma_{t-1},\varepsilon_{t-1},\textbf{s}_{\theta_{h}})$;
   11:       Update $\hat{Y}_{h,i}^{t-1}$ with Eq. (25);
@@
   13: End for
   14: Obtain the final image $\tilde{x}$;
   15: return $\tilde{x}$.
IV Experiment
IV-A Data Specification
@@
1) AAPM Challenge Data: The dataset used for training and testing is part of the AAPM Low-Dose Challenge[49] provided by the Mayo Clinic. In this study, 4,742 slices from 9 patients were used for training, while 12 slices from the remaining patient were used for testing. Each slice is 1 mm thick. To evaluate the performance of the proposed method with different sparse views, CT images were reconstructed using 60, 90, and 120 projections, and both qualitative and quantitative analyses were performed. Reference images are generated from 720 projection views using the FBP algorithm. For the fan-beam CT reconstruction task, ray-driven algorithms[50, 51] were employed to simulate the generation of projection data. The distance from the center of rotation to the source and detector is set at 40 cm and 40 cm, respectively. The detector is 41.3 cm wide, has 720 detector elements, and a total of 720 projection views are evenly distributed.
@@
2) CIRS Phantom Data: We use a dataset of high-quality CT scans to further analyze the performance of the proposed method, each with dimensions of 512×512×100 voxels, and a voxel dimension of 0.78×0.78×0.625 mm3. The data set was collected using the GE Discovery HD750 CT scanner combined with a bionic model provided by CIRS. The vacuum tube current is set to 600 mAs, the source-wheelbase separation of the CT system is 573mm, and the source-detector distance is 1010 mm. To evaluate the efficacy of the reconstruction algorithms across different sparse-view sampling scenarios, CT images were reconstructed using 60, 90, and 120 projection, the reconstruction results underwent both qualitative and quantitative analysis.
@@
3) Dental Arch Data: The data set is the clinical data collected by JIROX Dental CBCT device produced by YOFO (Hefei) Medical Technology Co., Ltd. The source-to-image distance (SID) of the device is 1700 mm, and the source-to-detector distance (SAD) is 1500 mm. The device operates at a voltage of 100 kV and a current of 6 mA. The detector array consisted of 768 × 768 elements, with each detector element measuring 0.2 × 0.2 mm. The dataset consists of 20 cases, and each case provides 200 slices. The sinograms have a size of 1536×1600, which correspond to 1600 uniformly-sampled projections within the angular range of $[0^{\circ},360^{\circ}]$. The dataset consists of full-view sinogram and their corresponding FBP reconstructions, with 512x512 image matrix. We randomly selected 1 case from the obtained image matrices and performed CT reconstruction using 60, 90, and 120 projections, respectively.
IV-B Implementation Details
In the experimental part of this study, we use Adam algorithm to train the model, and the learning rate is set to $10^{-3}$. The virtual mask are randomly designed in the following way. By simulating the sparse-view mask at 10, 20, 30, 60, 90, 120, and 180 projection. At the same time, use the Kaiming initialization to initialize the weights of the model. The method is implemented in Python using Operator Discretization Library (ODL)[52] and PyTorch on a personal workstation with a GPU card (Tesla V100-PCIE-16 GB). During the reconstruction process, the iteration parameter was set to $N=800$ for 60 views and $N=1200$ for 90 and 120 views. Our source code is publicly available on GitHub and can be accessed via the following link: https://github.com/yqx7150/SWARM. To evaluate the performance quantitatively, peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) and mean square error (MSE) were used.
TABLE I: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of AAPM Challenge Data Using Different Methods at 60, 90, 120 Views.
@@ IV-B Implementation Details
Views FBP[53] FBPConvNet[14] HDNet[18] GMSD[21] SWORD[22] SWARM
@@
Figure 3: Reconstruction images from 120 views using different methods with AAPM challenge data. (a) The reference image versus the images reconstructed by (b) FBP, (c) FBPConvNet, (d) HDNet, (e) GMSD, (f) SWORD, and (g) SWARM. The display window is [-50, 490] HU. The second line is the residual between the reference image and the reconstructed image.
@@
TABLE II: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of CIRS Phantom Data Using Different Methods at 60, 90, 120 Views.
Views FBP[53] FBPConvNet[14] HDNet[18] GMSD[21] SWORD[22] SWARM
@@
TABLE III: Reconstruction PSNR/SSIM/MSE$(10^{-3})$ of Dental Arch Data Using Different Methods at 60, 90, 120 Views.
Views FBP[53] FBPConvNet[14] HDNet[18] GMSD[21] SWORD[22] SWARM
@@
IV-C Reconstruction Experiments
1) AAPM Reconstruction Results: In order to assess the effectiveness of the proposed algorithm, we evaluated the performance and compared some representative methods including FBP[53], FBPConvNet[14], HDNet[18], GMSD[21], SWORD[22]. Sparse-view CT reconstructions were conducted using 60, 90, and 120 projection, respectively. Table I presents the PSNR, SSIM and MSE values for the reconstruction results of the AAPM challenge dataset. The best values for reconstructed images with different projection views are highlighted in bold. SWARM demonstrated superior performance over other reconstruction techniques in terms of mean values for the AAPM test data. In contrast, FBP and HDNet exhibited inferior numerical performance for CT reconstruction under sparse-view conditions. Furthermore, SWARM showed significant advantages over existing technologies.
Fig. 3 and Fig. 4 shows the visual reconstruction results of the AAPM test dataset. FBP exhibit suboptimal performance in image reconstruction, often resulting in severe blurring of fine details and pronounced edge artifacts. Although FBPConvNet has alleviated some of these limitations, it still suffers from insufficient edge sharpness and inadequate recovery of fine structures, leading to the omission of subtle image features. The dual-domain iterative model HDNet, which integrates information from both frequency and image domains, achieves certain improvements in reconstruction quality. However, it tends to introduce excessive smoothing, consequently causing the loss of critical structural information. Similarly, GMSD and SWORD inevitably compromise texture information during iterative reconstruction, thereby diminishing the realism and fidelity of reconstructed images. In contrast, the SWARM method demonstrates superior performance in preserving texture details and restoring edge contours. It maintains higher fidelity in representing structural features such as the shape and grayscale distribution of target regions, exhibiting significantly improved consistency with ground truth images. Both in terms of visual perception and quantitative metrics, SWARM consistently outperforms the aforementioned approaches.
2) CIRS Phantom Reconstruction Results: To further evaluate the generalization and robustness of the proposed unsupervised learning method, prior knowledge was learned from the AAPM challenge dataset, and the model performance was assessed using CIRS phantom data. Table II shows that SWARM significantly outperforms the other approaches in terms of quantitative metrics. The evaluation results of FBP and FBPConvNet on the CIRS phantom dataset are notably inferior, indicating considerable loss of image information. Although the quantitative metrics of HDNet and GMSD remain relatively stable, their reconstruction performance still shows a noticeable decline. SWORD shows relatively stable reconstruction performance in various indicators. However, it can still be observed that its reconstruction quality decreases to a certain extent in some test cases, revealing the limitation of insufficient generalization ability. In contrast, SWARM gets the best results, demonstrating its strong generalization and robustness.
As shown in Fig. 5, our method achieves excellent visual results. Images reconstructed using the conventional FBP method suffer from prominent streak artifacts and significant loss of structural details. FBPConvNet remains limited in restoring edge sharpness and texture details, resulting in blurred structural contours. Despite leveraging iterative optimization of dual-domain information, HDNet remains inadequate in accurately delineating internal structural boundaries. Similarly, GMSD and SWORD tend to over-smooth the reconstructed images, leading to the loss of texture information and blurring of fine structures. In contrast, SWARM demonstrates superior performance by effectively suppressing artifacts while preserving rich texture details and structural features. This results in a substantial enhancement of visual reconstruction quality, fully showcasing the method’s excellent generalization capability and robustness across diverse reconstruction scenarios.
3) Dental Arch Reconstruction Results: To evaluate the practicality of the proposed method, the model learned prior knowledge from the abdominal dataset of the AAPM challenge and assessed the model performance using the Panoramic Dental Arch data. Table III shows the significant advantages of SWARM in terms of quantitative metrics. Although other comparison methods remain relatively stable in the evaluation results or show slight improvements, they still fail to fully demonstrate their reconstruction capabilities. In contrast, SWARM shows obvious advantages with strong generalization ability and robustness, which provides an effective idea for solving the problem of SVCT reconstruction, and is expected to provide a potential reference for clinical diagnosis in stomatology in theory.
As shown in Fig. 6 and Fig. 7, our approach provides excellent visual results in the processing of oral data. In contrast, images reconstructed using FBP, FBPConvNet, HDNet, and GMSD exhibit significant streak artifacts and detail loss, particularly in the dental regions. While SWORD alleviates some of these issues, it still struggles with the restoration of structural details, especially in areas such as the mandibular molars, resulting in blurred images. In comparison, SWARM not only effectively suppresses streak artifacts but also achieves notable success in preserving texture details, particularly in the high-quality structural reconstruction of the mandibular molars. Overall, SWARM stands out in terms of visual reconstruction quality for oral images, providing clinicians with clearer and more accurate imaging support, thus enhancing the reliability and precision of diagnostic assistance.

### v3

@@ 3.3 Cascade Reconstruction of Sinogram and Wavelet
[…]
@@
where $\tilde{H}$$\cdot$] represents the full-frequency information, $\tilde{H}_{l}[\cdot]$ represents the low-frequency component and $\tilde{H}_{h}[\cdot]$ represents the high-frequency component. Through the iterative optimization mechanism of SRM and SHD, a smooth transition from global consistency to detail precision is effectively achieved.
3.4 Sinogram Generation
The sinogram generation stage enhances the global structural information in the image, thereby ensuring the consistency and integrity of the data during the processing. In this stage, it can be further refined into two interrelated subtasks, which respectively focus on the modeling of global information and the expression and optimization of structural features, so as to improve the expressive ability of the overall data and the reconstruction accuracy. This stage can be decomposed into the following two sub-problems:
@@ 3.4 Sinogram Generation
$\hat{y}_{s}^{t-1}=\underset{\hat{y}_{s}}{\arg\min}\ \|\hat{y}_{s}-\hat{y}_{s}^{t-\frac{1}{2}}\|^{2}_{2}+\nu_{s}R_{s}(\hat{y}_{s}).\vskip-5.0pt$ (16)
@@
$\hat{y}_{s}=(1-P(\wedge))\hat{y}_{s}+P(\wedge)y.\vskip-5.0pt$ (17)
@@
3.5 High-frequency generation in wavelet domain
After the preliminary reconstruction is completed in the projection domain during the first stage, the obtained sinogram is mapped to the wavelet domain. In this domain, the image is decomposed at multiple scales, from which the high-frequency components in three directions are extracted, and the low-frequency components are retained simultaneously, so as to fully capture the detailed structures and the overall contours in the image. The process can be described as follows:
@@ 3.5 High-frequency generation in wavelet domain
3.6 Domain transform stage
Finally, this process integrates the previously retained low-frequency information with the optimized high-frequency components. By fusing information across different frequency levels, a more comprehensive image representation is achieved. Subsequently, the combined data is mapped back to the projection domain through the inverse wavelet transform, as detailed below:
@@ 3.6 Domain transform stage
[…]
In summary, the iterative reconstruction stage of SWARM is shown in Fig. 2(c). In the actual reconstruction, through the iterative updates of the numerical SDE solver and the Langevin dynamics, high-quality full-view projection data is gradually obtained. Algorithm 1 provides a detailed description of both the training process and the reconstruction stage.
Algorithm 1 Training and Iterative Reconstruction Process.
@@
SRM Training: Sample a projected data $y$, apply a random mask $\hat{m}$ to generate $\tilde{Y}_{s}=y\odot\hat{m}$, and train the SRM network $\textbf{s}_{\theta_{s}}(\tilde{Y}_{s},t)$.
[…]
Iterative Reconstruction Process:
@@
  1: For t=T1t=T-1 to 0 do:
  2:    $\hat{Y}_{s}^{t-\frac{1}{2}}\leftarrow predictor(\hat{Y}_{s}^{t},\sigma_{t-1},\sigma_{t},\textbf{s}_{\theta_{s}})$;
  3:    Update $\hat{Y}_{s}^{t-\frac{1}{2}}$ with Eq. (17);
  4:    $\hat{Y}_{s}^{t-1}\leftarrow corrector(\hat{Y}_{s}^{t-\frac{1}{2}},\sigma_{t-1},\varepsilon_{t-1},\textbf{s}_{\theta_{s}})$;
  5:    Update $\hat{Y}_{s}^{t-1}$ with Eq. (17);
  6:    For each subband $i\in\{LH,HL,HH\}$ do:
  7:     $\hat{Y}_{h,i}^{t-\frac{1}{2}}\leftarrow predictor(\hat{Y}_{h,i}^{t},\sigma_{t-1},\sigma_{t},\textbf{s}_{\theta_{h}})$;
  8:     Update $\hat{Y}_{h,i}^{t-\frac{1}{2}}$ via Eq. (21);
  9:     $\hat{Y}_{h,i}^{t-1}\leftarrow corrector(\hat{Y}_{h,i}^{t-\frac{1}{2}},\sigma_{t-1},\varepsilon_{t-1},\textbf{s}_{\theta_{h}})$;
  10:     Update $\hat{Y}_{h,i}^{t-1}$ via Eq. (21);
  11:   End for
@@
  13: return $\tilde{x}$.
4 Experiment
4.1 Data Specification
4.1.1 AAPM Challenge Data
@@
The dataset used for training and testing is part of the AAPM Low-Dose Challenge[AAPM] provided by the Mayo Clinic. In this study, 9 patients were used for The distance from the center of rotation to the source and detector is set at 40 cm and 40 cm, respectively. The detector is 41.3 cm wide, has 720 detector elements, and a total of 720 projection views are evenly distributed.
4.1.2 CIRS Phantom Data
@@
We use a dataset of high-quality CT scans to further analyze the performance of the proposed method, each with dimensions of 512×512×100 voxels, and a voxel dimension of 0.78×0.78×0.625 mm3. The data set was collected using the GE Discovery HD750 CT scanner combined with a bionic model provided by CIRS. The vacuum tube current is set to 600 mAs, the source-wheelbase separation of the CT system is 573mm, and the source-detector distance is 1010 mm.
4.1.3 Dental Arch Data
@@
The data set is the clinical data collected by JIROX Dental CBCT device produced by YOFO (Hefei) Medical Technology Co., Ltd. The source-to-image distance (SID) of the device is 1700 mm, and the source-to-detector distance (SAD) is 1500 mm. The device operates at a voltage of 100 kV and a current of 6 mA. The detector array consisted of 768 × 768 elements, with each detector element measuring 0.2 × 0.2 mm. The dataset consists of 20 cases, and each case provides 200 slices. The dataset consists of full-view sinogram and their corresponding FBP reconstructions, with 512 × 512 image matrix. We randomly selected 1 case from the obtained image matrices.
4.2 Implementation Details
In our experiments, the model is trained using the Adam optimizer with a learning rate of $10^{-3}$ and Kaiming weight initialization. Model hyperparameters are set as follows: VE-SDE parameters $\sigma_{\min}=0.01$ and $\sigma_{\max}=378$. The noise step schedule $\epsilon_{t}$ follows the VE-SDE [song2020score], and the predictor–corrector (PC) step ratio is set to 1.
Sparse-view projections were simulated using inverse masks at 10, 20, 30, 60, 90, 120, and 180 views. The masks were generated under the sparse scanning model by uniformly random sampling without fixing random seeds, and independently applied to each sample. The 2D discrete wavelet transform (DWT) uses the Haar wavelet with two decomposition levels. Boundary handling is done via symmetric extension with matrix construction and cropping, ensuring stability and boundary preservation. High-frequency components in each sub-band are randomly selected with uniform probability.
In the fan-beam CT experiments, ray-driven algorithms [article1, article] simulated projection data, using ODL [rajpurkar2022ai]. The projection geometry includes 720 projection angles over $0,2\pi]$ and 720 detector elements over [360,360][-360,360] mm. Both the SOD and the SID are 500 mm, and the reconstruction domain is discretized on a $512\times 512$ grid. The reference images are generated from 720 projection views using the FBP algorithm. Our source code is publicly available on GitHub and can be accessed via the following link: https://github.com/yqx7150/SWARM. To evaluate the proposed method with different sparse views, CT images were reconstructed using 60, 90, and 120 projections. Performance was quantitatively assessed using PSNR, SSIM, and MSE.
[…]
@@ 4.2 Implementation Details
Views FBP[brenner2007computed] FBPConvNet[McCann2017Deep] HDNet[Hu2021Hybrid-Domain] GMSD[Guan2023Generative] SWORD[10403850] SWARM
@@
[…]
Views FBP[brenner2007computed] FBPConvNet[McCann2017Deep] HDNet[Hu2021Hybrid-Domain] GMSD[Guan2023Generative] SWORD[10403850] SWARM
@@
[…]
Views FBP[brenner2007computed] FBPConvNet[McCann2017Deep] HDNet[Hu2021Hybrid-Domain] GMSD[Guan2023Generative] SWORD[10403850] SWARM
@@
4.3 Reconstruction Experiments
4.3.1 AAPM Reconstruction Results
In order to assess the effectiveness of the proposed algorithm, we evaluated the performance and compared some representative methods including FBP[brenner2007computed], FBPConvNet[McCann2017Deep], HDNet[Hu2021Hybrid-Domain], GMSD[Guan2023Generative], SWORD[10403850]. Sparse-view CT reconstructions were conducted using 60, 90, and 120 projection, respectively. Table 1 presents the PSNR, SSIM and MSE values for the reconstruction results of the AAPM challenge dataset. The best values for reconstructed images with different projection views are highlighted in bold. SWARM indicating its superior performance in overall image quality and signal fidelity. The higher PSNR demonstrates that SWARM can more accurately restore the overall signal intensity of the reconstructed images, maintaining a high consistency with the ground truth. The optimal SSIM further reflects its enhanced capability in preserving structural information and local texture details. The substantial reduction in MSE further confirms the minimal reconstruction error, bringing the reconstructed images closer to the ground truth.
Fig. 3 shows the visual reconstruction results of the AAPM test dataset. FBP exhibits noticeable edge blurring and streak artifacts, with severe loss of fine structures. Although FBPConvNet alleviates some of these artifacts, edge sharpness remains insufficient, and certain local details are still not recovered. HDNet achieves a generally smoother reconstruction, but excessive smoothing leads to blurred key structural boundaries. Similarly, GMSD and SWORD suffer from texture loss and weakened detail preservation during iterative reconstruction. In contrast, SWARM more clearly restores the edges and structural details of lung tissue, with the reconstructed images showing higher consistency with the ground truth in terms of texture, morphology, and intensity distribution. In summary, SWARM not only preserves the structural integrity of the images but also enhances the overall signal accuracy, providing a more reliable imaging foundation for clinical applications.
4.3.2 CIRS Phantom Reconstruction Results
To further evaluate the proposed method, prior knowledge was learned from the AAPM challenge dataset, and the model performance was assessed using CIRS phantom data. Table 2 shows that SWARM significantly outperforms the other approaches in terms of quantitative metrics. The evaluation results of FBP and FBPConvNet on the CIRS phantom dataset are notably inferior, indicating considerable loss of image information. Although the quantitative metrics of HDNet and GMSD remain relatively stable, their reconstruction performance still shows a noticeable decline. SWORD shows relatively stable reconstruction performance in various indicators. However, it can still be observed that its reconstruction quality decreases to a certain extent in some test cases, revealing the limitation of insufficient generalization ability. In contrast, SWARM achieved the best reconstruction results, fully demonstrating its superior overall performance.
As shown in Fig. 4, FBP suffer from prominent streak artifacts and significant loss of structural details. FBPConvNet remains limited in restoring edge sharpness and texture details, resulting in blurred structural contours. Despite leveraging iterative optimization of dual-domain information, HDNet remains inadequate in accurately delineating internal structural boundaries. Similarly, GMSD and SWORD tend to over-smooth the reconstructed images, leading to the loss of texture information and blurring of fine structures. In contrast, SWARM demonstrates superior performance by effectively suppressing artifacts while preserving rich texture details and structural features. This results in a substantial enhancement of visual reconstruction quality, fully showcasing the method’s excellent generalization capability and robustness across diverse reconstruction scenarios.
4.3.3 Dental Arch Reconstruction Results
To evaluate the practicality of the proposed method, the model learned prior knowledge from the abdominal dataset of the AAPM challenge and assessed the model performance using the Dental Arch data. Table 3 shows the significant advantages of SWARM in terms of quantitative metrics. Although other comparison methods remain relatively stable in the evaluation results or show slight improvements, they still fail to fully demonstrate their reconstruction capabilities. In contrast, SWARM shows obvious advantages with strong generalization ability and robustness, which provides an effective idea for solving the problem of SVCT reconstruction, and is expected to provide a potential reference for clinical diagnosis in stomatology in theory.
As shown in Fig. 5 and Fig. 6, SWARM demonstrates excellent structural fidelity in oral image reconstruction. In contrast, FBP, FBPConvNet, HDNet, and GMSD exhibit noticeable streak artifacts and loss of detail in the dental region, particularly around the mandibular molars, resulting in unclear tooth morphology and indistinct boundaries of periodontal tissues, which may affect clinical assessment of tooth structures and lesions. Although SWORD partially alleviates these artifacts, it still produces blurring in regions with complex dental structures, making fine features such as interproximal spaces difficult to fully resolve. By comparison, SWARM effectively suppresses streak artifacts while preserving clear boundaries and rich textural details of teeth and surrounding tissues, with particularly outstanding performance in reconstructing the mandibular molar structures. The images reconstructed by this method provide clear visualization of dental regions, offering intuitive and discernible information for clinical dentistry, thereby enhancing diagnostic accuracy and reliability in treatment planning.